Please explain how you simplify fractions broken down and what is 15/90 in simplest form? How did u get it and how do u turn the answer into a decimal?
Found 2 solutions by rfer, drk:
Answer by rfer(16322)

15/90=1/6=0.167
15 divides into 90 6 times
1 divided by 6 gives the decimal

Answer by drk(1908)

In general to simplify or reduce fractions, you need to find a common factor or apply prime factorization to the numbers. We keep simplifying until there are no more common factors.
For example, 15/90
15 divided by 5 = 3
90 divided by 5 = 18
So, we go from 15/90 to 3/18
--
3 divided by 3 = 1
18 divided by 3 = 6
So, we go from 3/18 = 1/6.
--
To go from fraction to decimal, divide bottom into top; 1 inside and 6 outside division bar
We get .1666666